OpenStudy (anonymous):

Use only the constant rule and exponent rule.. show your solution. I know how to use the quotient rule but our class did not discuss it yet so i need to stick with two rules.. Differentiate:

OpenStudy (anonymous):

\[s(t)=\frac{ t ^{5}-3t ^{2} }{ 2t}, t>0\]

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

Do you know the product rule, though?

OpenStudy (anonymous):

Nevermind, you don't need it.

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @Data_LG2 \[s(t)=\frac{ t ^{5}-3t ^{2} }{ 2t}, t>0\] \(\color{blue}{\text{End of Quote}}\) Since \(t\not=0\), you can divide: \[s(t)=\frac{1}{2}\left(t^4-3t\right)\]

OpenStudy (anonymous):

ahh.. got it thanks!!